On January 1, 2018, Wirth Corporation, a publicly traded company, had these shareholders equity accounts: Common shares (unlimited number of shares authorized, 220,000 shares issued).........$2,200,000 Retained earnings..............................................................................................................1,080,000 Accumulated other comprehensive income..........................................................................120,000 During the year, the following transactions occurred: Jan. 15 Declared a $1 per share cash dividend to shareholders of record on January 31, payable February 15. Apr. 16 Declared a 10% stock dividend to shareholders of record on April 30, distributable May 16. On April 16, April 30, and May 15, the share prices were $15, $13.50, and $14, respectively. Oct. 1 Effected a 2-for-1 stock split. On October 1, the share price was $20. Dec. 31 Determined that net income for the year was $700,000. Instructions (a) Record the above transactions, including any required entries to close dividends and net income. (b) Open T accounts as required and post to the shareholders equity accounts. (c) Prepare a statement of changes in equity for the year. (d) Prepare the shareholders equity section of the statement of financial position at December 31.
Answer #1
(a) Date Account titles and explanation Debit Credit
Jan 15. Dividends (220000*1) 220000
Dividend payable 220000
(Dividend declared)
Jan 31. No entry
Feb 15. Dividend payable 220000
Cash 220000
(Dividend paid)
Apr 16. Stock dividend (220000*10%*15) 330000
Common shares distributable (220000*10%*10) 220000
Paid-in capital in excess of par-common stock 110000
(Stock dividend declared)
Apr 30. No entry
May 15. Common shares distributable 220000
Common shares 220000
(Stock dividend distributed)
Oct 1. No entry
Dec 31. Income summary 700000
Retained earnings 700000
(Net income recorded)
Dec 31. Retained earnings 550000
Dividends 220000
Stock dividend 330000
(Dividend closed)
(b) Common shares Paid-in capital in excess of par-common stock
Date Debit Date Credit Date Debit Date Credit
Jan 1. 2200000 Jan 1. 0
May 15. 220000 Apr 16. 110000
0 2420000 0 110000
Bal. 2420000 Bal. 110000
Retained earnings Accumulated other comprehensive income
Date Debit Date Credit Date Debit Date Credit
Dec 31. 550000 Jan 1. 1080000 Jan 1. 120000
Dec 31. 700000
550000 1780000 0 120000
Bal. 1230000 Bal. 120000
c) Statement of changes in equity
Common shares Paid-in capital in excess of par-common stock Retained earnings Accumulated other comprehensive income
Beginning balance 2200000 0 1080000 120000
Cash dividend paid -220000
Stock dividend paid 220000 110000 -330000
Net income 700000
Ending balance 2420000 110000 1230000 120000
d) Shareholder's equity
$
Common shares
(Unlimited number of shares authorized,484000
shares issued,par value $ 5) 2420000
Paid-in capital in excess of par-common stock 110000
Retained earnings 1230000
Accumulated other comprehensive income 120000
Total stockholder's equity 3880000
Note:
Par value for common shares at the beginning=2200000/220000=$10 per share
Stock split of 2:1
New par value=10*(1/2)=$ 5 per share
Number of common shares before stock split=220000+22000=$ 242000
Number of common shares after stock split=242000*(2/1)=484000 shares